# inkscapelatex=false

### v1

@@ Abstract
Skeleton-based action recognition (SAR) in videos is an important but challenging task in computer vision. The recent state-of-the-art models for SAR are primarily based on graph convolutional neural networks (GCNs), which are powerful in extracting the spatial information of skeleton data. However, it is yet clear that such GCN-based models can effectively capture the temporal dynamics of human action sequences. To this end, we propose the DevLSTM module, which exploits the path development – a principled and parsimonious representation for sequential data by leveraging the Lie group structure. The path development, originated from Rough path theory, can effectively capture the order of events in high-dimensional stream data with massive dimension reduction and consequently enhance the LSTM module substantially. Our proposed G-DevLSTM module can be conveniently plugged into the temporal graph, complementing existing advanced GCN-based models. Our empirical studies on the NTU60, NTU120 and Chalearn2013 datasets demonstrate that our proposed hybrid model significantly outperforms the current best-performing methods in SAR tasks. The code is available at https://github.com/DeepIntoStreams/GCN-DevLSTM.
@@ 1 Introduction
Action recognition has a wide range of applications in diverse fields such as human-computer interaction, performance assessment in athletics, virtual reality, and video monitoring Liu et al. (2019a). In comparison to video-based action recognition, skeleton-based action recognition (SAR) exhibits better robustness to the variety of light conditions and its efficiency in computation and storage Chen et al. (2021); Shi et al. (2019) while preserving data privacy. Skeleton data can be obtained either by localisation of 2D/3D human joints using the depth camera Liu et al. (2019a) or pose estimation algorithms from the videos Cao et al. (2017).
@@
Despite extensive literature devoted to SAR, the challenge still remains open. The key step towards accurate SAR models lies in designing an effective spatio-temporal representation of the skeleton sequence. Previously, Recurrent Neural Network (RNN) was commonly used due to its strength in capturing the temporal dynamics. However, since STGCN Yan et al. (2018) was proposed to capture the spatio-temporal features, which contain spatial graph convolution and temporal convolution, graph convolutional networks (GCNs) have taken precedence. The majority of subsequent GCN-based approaches Chen et al. (2021); Shi et al. (2019, 2020); Cheng et al. (2020b, a); Song et al. (2020) have largely followed the STGCN framework, while they mainly focus on the improvement of the spatial graph convolution, with minimal emphasis on improving temporal modules. All of these methods rely on $K\times 1$ convolution operation to capture temporal features, where $K$ represents the kernel size of the convolutional operation in the temporal receptive field. However, whether convolution operation is most effective for capturing temporal dynamics is in question.
@@
On the other hand, Rough path theory  Lyons (1998) originated as a branch of stochastic analysis to make sense of controlled differential equations driven by highly oscillatory paths. More recently, the applications of rough path theory in sequential data analysis are emerging. In particular, the path signature, as a fundamental object of Rough path theory, has successful applications in various domains ranging from finance to healthcare, demonstrating its capacity as a principled and efficient feature of time series. By embedding time series into the continuous path space, the signature offers a unified treatment to handle variable length and irregular sampling. Signature-based algorithms show promise across diverse fields, including SAR task Liao et al. (2021); Yang et al. (2022); Ahmad et al. (2019); Li et al. (2019a); Cheng et al. (2023). However, the path signature may face challenges such as the curse of dimensionality caused by high path dimension and lack of adaptability due to its deterministic nature Lou et al. (2022). To address these, Lou et al. Lou et al. (2022) propose the path development from rough path theory, as a trainable alternative to path signature. The path development enjoys the desirable properties of the signature while reducing feature dimension significantly.  Lou et al. (2022) shows that combining the path development under the appropriate Lie group with LSTM Hochreiter and Schmidhuber (1997) successfully alleviates the gradient vanishing issues of LSTMs, leading to significant enhancement across various sequential tasks.
@@
This paper places a primary emphasis on enhancing the temporal module within our framework. While retaining the effectiveness of graph convolution for modelling spatial relationships among joints, we introduce a novel module, G-DevLSTM, for handling temporal relationships of the graph with the node feature as time series. More specifically, we adopt a topology-non-shared graph from CTRGCN Chen et al. (2021) where each channel has an adjacency matrix while in contrast to employing only convolutional operation in the temporal module, we also integrate path development and LSTM in our framework. This combination enables our temporal block to capture more discriminative temporal features, leading to model performance improvement. Experimental results on three commonly used datasets show that our approach outperforms other methods, demonstrating the effectiveness of our proposed temporal module.
We summarise our main contributions as follows:
@@
  • We design a novel and principled graph development (G-Dev) layer for analysing temporal graph data by extending the path development originally applied to vector-valued time series. As a generic network on graph data with time-varying node attributes, G-Dev exhibits efficacy and robustness to irregular sampling and missing data.
@@ 2.1 Skeleton-based Action Recognition
SAR is a task where we need to consider both spatial and temporal relationships. Earlier methods normally use convolutional neural networks (CNNs) Chéron et al. (2015); Liu et al. (2017b) and RNN Du et al. (2015); Liu et al. (2017a) to address both spatial and temporal challenges without considering structural topology. STGCN Yan et al. (2018) is the first method to capture spatial and temporal relationships using GCN, taking the skeleton topology of the graph into consideration while this topology is not learnable, with only the spatial connection between adjacent joints being connected. As a result, some actions required to consider non-adjacent joints may perform poorly. Liu et al. (2020) addresses this by introducing a learnable adjacency matrix alongside the original manually defined one, enabling connections between non-neighboring joints. Yet, this approach maintains the same graph for all action inputs, which is not ideal, as different action classes may benefit from different graph topologies. To address this, Li et al. (2019b); Shi et al. (2019, 2020) propose adaptive graph convolutional network, allowing the topology to be learnable and data dependent while they still use the same topology for all channels. Cheng et al. (2020a); Chen et al. (2021) further propose non-shared topology which uses distinct topology for each channel. Current GCN-based methods put much effort into the enhancement of the graph, with limited attention to temporal design. Most GCN-based methods mainly rely on CNNs to capture temporal dynamics. Although studies such as Si et al. (2019); Zhao et al. (2019); Liao et al. (2021) utilise LSTM to capture temporal features while also using GCN to build spatial correlations, these models unfortunately experience a performance lag compared to those leveraging both GCN and CNN.
@@ 2.2 Path signature & Path Development
Path signature, introduced by Chen Chen (1958), is a representation of a piecewise regular path. This representation involves a collection of iterated integrals, which faithfully represent the original path. Lyons further broadens the scope from paths with bounded variation Hambly and Lyons (2010) to paths with finite p-variation for any $p\geq 1$ Boedihardjo et al. (2016). There are plenty of applications of path signature in various machine learning areas, containing handwritten recognition Xie et al. (2017), writer identification Yang et al. (2015), financial data analysis Lyons et al. (2014), time-series data generation Ni et al. (2021) and SAR Liao et al. (2021); Cheng et al. (2023); Yang et al. (2022); Ahmad et al. (2019); Li et al. (2019a). In the specific context of SAR, Ahmad et al. (2019) focuses on employing path signature solely for extracting spatial features while Yang et al. (2022); Li et al. (2019a) utilise path signature to obtain both temporal-spatial features. However, relying solely on path signature to capture spatial relationships may be suboptimal as it is not data adaptive and only considers local spatial correlations. Recognising the efficacy of GCN in modelling spatial relationships, Liao et al. (2021) leverages GCN to extract spatial features and employ path signature to capture temporal dynamics. Unfortunately, the network tends to overfit rapidly when increasing its depth due to the excessively large output dimension of the signature.
Trainable path development, as presented in the work Lou et al. (2022) is based on the development of a path à la Cartan Driver (1995), a significant mathematical tool in rough path theory, particularly for the examination of the inverse of the signature Lyons and Xu (2017) and the uniqueness of expected signature Hambly and Lyons (2010); Chevyrev and Lyons (2016). Chevyrev and Lyons (2016) demonstrates that given an appropriate Lie algebra, the path development has universal and characteristic properties, just like path signature, suggesting the potential of path development as a promising feature representation for sequential data. Moreover, path development has been successfully applied in PCF-GAN Lou et al. (2023) to generate high-quality time series data due to the favourable non-commutativity and group structure of the unitary feature.
This paper is built upon the path development. Instead of solely relying on CNN to capture temporal dynamics, we integrate path development with LSTM and propose a new temporal module to capture more discriminate temporal features.
@@ 3.1 The development of a Path
Thanks to the multiplicative property of the path development (Lemma 2.4 in Lou et al. (2022)), the development of two paths $X\in V([0,s],\mathbb{R}^{d})$ and $Y\in V([s,t],\mathbb{R}^{d})$ under $M$ is the same as the development of the concatenation of these two path, in formula,
@@
The path development enjoys several desirable theoretical properties, which makes itself a principled and efficient representation of sequential data. We summarise the properties, which are directly relevant to action recognition as follows.
Invariance under Time-reparametrisation
Similar to the path signature, the path development is invariant under time reparametrisation. This property ensures that the development of a path remains unchanged irrespective of the speed at which the path is traversed. It is well suited to action recognition, as the prediction of actions should not depend on the speed of the actions or the frame rate. Time-reparametrisation invariance of path development may bring massive dimension reduction and enhance the robustness to variation in speed.
Uniqueness of the path development
Intuitively, the path development effectively capture the order of events in time series, leveraging the non-commutative nature of matrix multiplication. Changing the order of events may lead to different development in general, highlighting its capacity to capture the temporal dynamics of sequential data. It is proven in Chevyrev and Lyons (2016) the path development is a characteristic and universal feature of an un-parameterised path in $V_{1}([0,T],\mathbb{R}^{d})$ under the appropriate Lie group. In other words, transforming a path of finite length to its development map $M\mapsto\text{Dev}_{M}(X)$ can uniquely determine the path $X$ and yield no information loss.
@@ 3.2 Path Development Layer
Proposed in  Lou et al. (2022), the path development layer is a novel neural network inspired by the path development by exploiting the representation of the matrix Lie group. More specifically, this layer utilises the parameterised linear map $M\in L(\mathbb{R}^{d},\mathfrak{g})$ by its linear coefficients $\theta\in\mathfrak{g}^{d}$.
@@
The path development layer, denoted by $D_{\theta}$, is defined as a mapping $\mathbb{R}^{d\times(N+1)}\rightarrow G$, which transforms any input time series $x=(x_{1},x_{2},...,x_{N})$ to the development $z_{N}\in G$ under the trainable linear map $M_{\theta}\in L(\mathbb{R}^{d},\mathfrak{g})$. Here for $n\in\{0,...,N-1\}$, we set
@@
Dimension reduction
The output of the path development has dimension $m^{2}$, where $m$ is the matrix size. Note that its dimension does not depend on the path dimension $d$ and time dimension $T$, resulting in the massive dimension reduction. This is in stark contrast to the signature feature of degree $k$, whose dimension $\left(\sum_{i=0}^{k}{d^{i}}=\frac{d^{k+1}-1}{d-1}\right)$ grows geometrically w.r.t $d$. Moreover, the path development is well suited to extract information from high-frequency time series or long time series without concern on the curse of dimensionality.
Robustness to irregular sampling
As the path development can be interpreted as a solution to the differential equation (Eq. (1), it is a continuous time series model. Similar to other continuous time models such as signature and Neural CDEs Kidger et al. (2020), the path development exhibits robustness in handling time series data with irregular sampling.
Figure 1: The work flow of G-Dev layer.
Figure 2: (a) The pipeline of our proposed approach consists of $N$ blocks, with each block containing a GCN module and a DevLSTM module. (b) The detail of the DevLSTM module of (a) is provided here. (c) The flowchart illustrates the sliding window setting for the path development layer in (b).
@@ 3.3 G-Dev Layer: Apply path development layer to a temporal graph
In this subsection, we propose the so-called G-Dev Layer, which serves as a generalisation of the development layer from the $d$-dimensional time series to the temporal graph. Throughout this paper, the temporal graph $\mathfrak{G}=(\mathcal{V},\mathcal{E},X)$ refers to the graph with each node associated with $d$-dimensional time series features. Here $\mathcal{V}:=\{v^{1},\cdots,v^{M}\}$ denotes the vertex set and $\mathcal{E}$ can be represented by the adjacent matrix $A\in\mathbb{R}^{m\times m}$. Let $X^{i}=(x_{t_{1}},\cdots,x_{t_{N_{i}}})\in\mathbb{R}^{d\times N_{i}}$ denote the features of the $i^{th}$ vertex, where $d$ and $N_{i}$ are the path and time dimension, respectively. The time dimension $N_{i}$ may vary across vertices $v_{i}$ for $i\in\{1,\cdots,M\}$. The feature sets $X=(X^{(1)},\cdots,X^{(M)})$ is obtained by stacking $X^{i}$ for all $i\in\{1,\cdots,M\}$. Clearly, the skeleton sequence is an example of temporal graph data.
The proposed G-Dev Layer is the transformation of the temporal graph by applying the development layer to the time-augmented transformation of each feature vector $X_{i}$ of the vertex $v_{i}$. The G-Dev layer is a trainable layer with model parameters $\theta\in\mathfrak{g}^{d}\subset\mathbb{R}^{m\times m\times d}$. As shown in Figure  1, at each vertex $V_{i}$, we apply the development transformation to its associated time series feature $X^{(i)}$, resulting in the matrix-valued output $Z_{t_{N}}\in G\subset\mathbb{R}^{m\times m}$.

### v2

# GCN-DevLSTM: Path Development for Skeleton-Based Action Recognition

@@ Abstract
Skeleton-based action recognition (SAR) in videos is an important but challenging task in computer vision. The recent state-of-the-art (SOTA) models for SAR are primarily based on graph convolutional neural networks (GCNs), which are powerful in extracting the spatial information of skeleton data. However, it is yet clear that such GCN-based models can effectively capture the temporal dynamics of human action sequences. To this end, we propose the G-Dev layer, which exploits the path development – a principled and parsimonious representation for sequential data by leveraging the Lie group structure. By integrating the G-Dev layer, the hybrid G-DevLSTM module enhances the traditional LSTM to reduce the time dimension while retaining high-frequency information. It can be conveniently applied to any temporal graph data, complementing existing advanced GCN-based models. Our empirical studies on the NTU60, NTU120 and Chalearn2013 datasets demonstrate that our proposed GCN-DevLSTM network consistently improves the strong GCN baseline models and achieves SOTA results with superior robustness in SAR tasks. The code is available at https://github.com/DeepIntoStreams/GCN-DevLSTM.
@@ 1 Introduction
Action recognition has a wide range of applications in diverse fields such as human-computer interaction, performance assessment in athletics, virtual reality, and video monitoring Liu et al. (2019a). In comparison to video-based action recognition, skeleton-based action recognition (SAR) exhibits better robustness to the variety of light conditions and its efficiency in computation and storage Chen et al. (2021); Shi et al. (2019) while preserving data privacy. Skeleton data can be obtained either by localization of 2D/3D human joints using the depth camera Liu et al. (2019a) or pose estimation algorithms from the videos Cao et al. (2017).
@@
Despite extensive literature devoted to SAR, the challenge still remains open. The key step towards accurate SAR models lies in designing an effective spatio-temporal representation of the skeleton sequence. Previously, Recurrent Neural Networks (RNNs) were commonly used due to their strength in capturing temporal dynamics. The advent of Spatial Temporal Graph Convolutional Networks (STGCN) has shifted the focus towards Graph Convolutional Networks (GCNs), which incorporate both spatial graph convolution and temporal convolution. The majority of subsequent GCN-based approaches Chen et al. (2021); Shi et al. (2020); Cheng et al. (2020b, a); Song et al. (2020); Zhou et al. (2022), following the STGCN framework, focus on the improvement of the spatial graph convolution, with minimal emphasis on improving temporal modules. These methods utilize 1-D temporal convolution operation to capture temporal features. However, whether convolution operation is most effective for capturing temporal dynamics is in question.
Figure 1: The work flow of G-Dev layer.
@@
On the other hand, Rough path theory Lyons (1998) originated as a branch of stochastic analysis to make sense of controlled differential equations driven by highly oscillatory paths. More recently, the applications of rough path theory in sequential data analysis are emerging. In particular, the path signature, as a fundamental object of Rough path theory, has successful applications in various domains ranging from finance to healthcare, demonstrating its capacity as a principled and efficient feature of time series. By embedding time series into the continuous path space, the signature offers a unified treatment to handle variable length and irregular sampling. Signature-based algorithms show promise across diverse fields, including SAR task Liao et al. (2021); Yang et al. (2022); Ahmad et al. (2019); Li et al. (2019a); Cheng et al. (2023). However, the path signature may face challenges such as the curse of dimensionality caused by high path dimension and lack of adaptability due to its deterministic nature Lou et al. (2022). To address these, Lou et al. Lou et al. (2022) propose the path development from rough path theory, as a trainable alternative to path signature. The path development enjoys the desirable properties of the signature while reducing feature dimension significantly. Lou et al. (2022) shows that combining the path development under the appropriate Lie group with LSTM Hochreiter and Schmidhuber (1997) alleviates the gradient vanishing of LSTMs, leading to significant enhancement across various sequential tasks.
@@
This paper places a primary emphasis on enhancing the temporal module within our framework. To this end, we introduce a novel G-Dev layer (as shown in Figure 1) on the temporal graph, which refers to the graph with the node feature as time series. This layer utilizes the sliding window to summarize the local information via the principled path development, offering time dimension benefit and robustness to missing data and irregular sampling. Moreover, we propose the GCN-DevLSTM network, composed of the CNN-based branch and G-Devlopment-based branch for SAR. While retaining the effectiveness of graph convolution for modeling spatial relationships among joints, the GCN-DevLSTM network exploits the hybrid capacity of G-Dev layer and LSTM for handling temporal dynamics. Our GCN-DevLSTM is flexible to incorporate with different GCN backbones. This combination captures discriminative temporal features and improves performance. Experimental results on three commonly used datasets Chalearn 2013, NTU 60 and NTU 120 show that our approach with the GCN backbone being topology-non-shared graph from CTRGCN Chen et al. (2021) outperforms other methods, demonstrating the effectiveness and robustness of our proposed temporal module.
We summarize our main contributions as follows:
@@
  • We design a novel and principled graph development (G-Dev) layer for analyzing temporal graph data by extending the path development originally applied to vector-valued time series. As a generic network on graph data with time-varying node attributes, G-Dev exhibits efficacy and robustness to irregular sampling and missing data.
@@ 2.1 Skeleton-based Action Recognition
The main challenge in SAR is extracting the effective representation that captures both spatial and temporal dependency of skeleton sequences. Earlier methods typically use convolutional neural networks (CNNs) Chéron et al. (2015); Liu et al. (2017b) and RNN Du et al. (2015); Liu et al. (2017a) to address this challenge without considering structural topology. STGCN Yan et al. (2018) is the first method to capture spatial and temporal relationships using GCN, taking the skeleton topology of the graph into consideration while this topology is not learnable, with only the spatial connection between adjacent joints being connected. As a result, some actions that rely on non-adjacent joints may lead to poor results. Liu et al. (2020) tackles this issue by introducing a learnable adjacency matrix alongside the original manually defined one, enabling connections between non-neighboring joints. Yet, this approach maintains the same graph for all action inputs, which might not be ideal, as different action classes may benefit from different graph topologies. To address this, Li et al. (2019b); Shi et al. (2019, 2020) propose adaptive GCN with the learnable and data-driven topology that shares among all channels. Cheng et al. (2020a); Chen et al. (2021) further propose the non-shared topology for each channel. Current GCN-based methods mainly focus on the enhancement of the graph representation, with limited attention to temporal design. In particular, most GCN-based methods typically use CNNs to capture temporal dynamics. Although studies such as Si et al. (2019); Zhao et al. (2019); Liao et al. (2021) attempt to incorporate LSTM with GCN for the SAR tasks, these methods still significantly underperform the sota GCN models that combine both GCN and CNN on large-scale action datasets.
@@ 2.2 Path signature & Path Development
The path signature is a core object of rough path theory, which provides a rigorous way of making sense of the solution to controlled differential equations driven by rough signals Lyons et al. (2007); Lyons (1998). Playing a role as the non-commutative monomials on the path space, the path signature can be used as a faithful feature of time series Lyons (2014); Levin et al. (2013), which leads to applications for various sequential data learning tasks, such as handwritten recognition Xie et al. (2017), writer identification Yang et al. (2015), financial data analysis Lyons et al. (2014), time-series data generation Ni et al. (2021) and SAR Liao et al. (2021); Cheng et al. (2023); Yang et al. (2022); Ahmad et al. (2019); Li et al. (2019a). In the specific context of SAR, Ahmad et al. (2019) focuses on employing path signature solely for extracting spatial features while Yang et al. (2022); Li et al. (2019a) utilize path signature to obtain both temporal and spatial features. However, relying solely on path signature to capture spatial relationships may be suboptimal, as it is not data adaptive and only considers local spatial correlations. Liao et al. (2021) leverages GCN to extract spatial features and employ the log signature to capture temporal dynamics. However, the network tends to overfit rapidly when increasing its depth due to the curse of dimensionality of the signature for the high dimensional path.
More recently, the path development layer Lou et al. (2022) is proposed as a general sequential layer on time series data as a trainable alternative to the signature feature. It has useful applications in supervised learning on sequential data Lou et al. (2022) and synthetic time generation lou2024pcf. This layer is inspired by the (Cantan) development of a path Driver (1995). The path development serves as a useful mathematical tool in rough path theory for proving the properties of the path signature (see e.g.,  Hambly and Lyons (2010); Chevyrev and Lyons (2016)). In particular,Chevyrev and Lyons (2016) proves that under an appropriate Lie group, the path development is a universal and characteristic feature of the path signature, which provides the theoretical support for the development layer Lou et al. (2022). Building upon Lou et al. (2022), our work extends the application of the path development from the vector-valued time series to that of the temporal graph.
@@ 3.1 The development of a Path
Thanks to the multiplicative property of the path development (Lemma 2.4 in Lou et al. (2022)), the development of two paths $X\in V([0,s],\mathbb{R}^{d})$ and $Y\in V([s,t],\mathbb{R}^{d})$ under $M$ is the same as the development of the concatenation of these two paths, in formula,
@@
The path development enjoys several desirable theoretical properties, making it a principled and efficient representation of sequential data. We summarize the following two properties, that are directly relevant to action recognition.
Invariance under Time-reparametrization Similar to the path signature, the path development is invariant under time reparametrization. This property ensures that the development of a path remains unchanged irrespective of the speed at which the path is traversed. It is well suited to action recognition, as the prediction of actions should not depend on the speed of the actions or the frame rate. Time-reparametrization invariance of path development may bring massive dimension reduction and enhance the robustness to variation in speed.
Uniqueness of the path development Intuitively, the path development effectively captures the order of events in time series, leveraging the non-commutative nature of matrix multiplication. Changing the order of events may lead to different development in general, highlighting its capacity to capture the temporal dynamics of sequential data. It is proven in Chevyrev and Lyons (2016) the path development is a characteristic and universal feature of an un-parameterized path in $V_{1}([0,T],\mathbb{R}^{d})$ under the appropriate Lie group. In other words, transforming a path of finite length to its development map $M\mapsto\text{Dev}_{M}(X)$ can uniquely determine the path $X$ up to translation. Thus, the path development combined with the starting point effectively summarizes the path without any loss of information.
@@ 3.2 Path Development Layer
Proposed in  Lou et al. (2022), the path development layer is a novel neural network inspired by the path development by exploiting the representation of the matrix Lie group. More specifically, this layer utilizes the parameterized linear map $M\in L(\mathbb{R}^{d},\mathfrak{g})$ by its linear coefficients $\theta\in\mathfrak{g}^{d}$.
@@
The path development layer, denoted by $D_{\theta}$, is defined as a mapping $\mathbb{R}^{d\times(N+1)}\rightarrow G$, which transforms any input time series $x=(x_{1},x_{2},...,x_{N})$ to the development $z_{N}\in G$ under the trainable linear map $M_{\theta}\in L(\mathbb{R}^{d},\mathfrak{g})$. For $n\in\{0,...,N-1\}$, we set
@@
Dimension reduction The output of the path development has dimension $m^{2}$, where $m$ is the matrix size. Note that its dimension does not depend on the path dimension $d$ and time dimension $T$, resulting in the massive dimension reduction. This is in stark contrast to the signature feature of degree $k$, whose dimension $\left(\sum_{i=0}^{k}{d^{i}}=\frac{d^{k+1}-1}{d-1}\right)$ grows geometrically w.r.t $d$. Moreover, the path development is well suited to extract information from high-frequency time series or long time series without concern on the curse of dimensionality.
Robustness to irregular sampling As the path development can be interpreted as a solution to the differential equation (Eq. (1), it is a continuous time series model. Similar to other continuous time models such as signature and Neural CDEs Kidger et al. (2020), the path development exhibits robustness in handling time series data with irregular sampling.
Computational Efficiency of analyzing online data Multiplicative property of the path development allows incremental computation of the development layer in streaming data, thereby reducing both computational load and memory requirements.
@@ 3.3 G-Dev Layer: Apply path development layer to a temporal graph
In this subsection, we propose the so-called G-Dev Sequence Layer, which extracts the local information of the temporal graph using the development layer in a rolling window fashion. The core component of the G-Dev Sequence layer serves as a generalization of the development layer from the $d$-dimensional time series to the temporal graph.
Throughout this paper, the temporal graph $\mathfrak{G}_{X}=(\mathcal{V},\mathcal{E},X)$ refers to the graph with each node associated with $d$-dimensional time series features $X$. Here $\mathcal{V}:=\{v^{1},\cdots,v^{M}\}$ denotes the vertex set and $\mathcal{E}$ can be represented by the adjacent matrix $A\in\mathbb{R}^{M\times M}$. We consider the general case where the time dimension of each node may vary across the nodes. More specifically, Let $X^{i}=(x_{t_{1}},\cdots,x_{t_{L_{i}}})\in\mathbb{R}^{d\times L_{i}}$ denote the features of the $i^{th}$ vertex, where $d$ and $N_{i}$ are the path and time dimension, respectively. The feature sets $X=(X^{(1)},\cdots,X^{(M)})$ is obtained by stacking $X^{i}$ for all $i\in\{1,\cdots,M\}$. Clearly, the skeleton sequence is an example of temporal graph data.
G-Dev module. Let us first define the G-Dev module, which is a trainable layer with model parameters $\theta\in\mathfrak{g}^{d}\subset\mathbb{R}^{m^{2}\times d}$, where $m$ is the matrix order of development. It maps a temporal graph $\mathfrak{G}_{X}=(\mathcal{V},\mathcal{E},X)$ to a static graph with matrix-valued features $\mathfrak{G}_{Z}=(\mathcal{V},\mathcal{E},Z)$ by applying the path development to the time-augmented transformation of each feature vector $X_{i}$ of the vertex $v_{i}$, as shown in Figure  1. In formula, for each vertex index $i\in\{1,\cdots,M\}$, the output feature is $Z=(Z^{(i)})_{i=1}^{M}$, where $Z^{(i)}=D_{\theta}(X^{(i)})\in G\subset\mathbb{R}^{m\times m}$.
Built on top of the G-Dev module, we employ a sliding window strategy to construct the following G-Dev Sequence layer, which can summarize (potentially long) time series over coarse time partitions to reduce the time dimension effectively while retaining the fine-grained information by using the low-dimensional representation.
Without loss of generality, let us focus on the case where the graph features $X$ share the same time stamps $(t_{i})_{i=0}^{L-1}$ across vertices [^1] Otherwise, we apply linear interpolation for data imputation, which would not affect the development feature thanks to invariance to time re-parameterization.. Fix the window (kernel) size $K$ and stride $S$. Let $L^{\prime}=\lceil\frac{L-k}{S}\rceil$. By applying padding to $Y$, for any $Y\in\mathbb{R}^{d\times N}$, we define a sequence of time series $\tilde{Y}_{j}$:
$\tilde{Y}_{j}=(Y_{j\times S},\cdots Y_{j\times S+K-1})\in\mathbb{R}^{d\times K},\forall j\in\{0,\cdots,L^{\prime}-1\}.$ (5)
Definition 3.3 (G-Dev Sequence layer).
[…]
The G-Dev Sequence layer (Definition 3.3) can be easily extended to the moving window with the dilation parameter $D$, where we collect the path information at every $D$ time steps during each moving window. To guarantee no information loss of the G-Dev sequence layer, it is crucial to choose $S$ and $K$ to ensure a minimum one-point overlap between two consecutive moving windows.
Figure 2: (a) The pipeline of our proposed approach consists of $N$ blocks, with each block containing a GCN module and a DevLSTM module. (b) The detail of the DevLSTM module.